The local Mars grocery store uses a periodic review model to stock their paper towels. After collecting some data, the firm was able to establish that the daily demand for paper towels is 17 packs. The store reviews its inventory once a week and places an order on Sunday. The lead time to get the product from the local distribution center is 4 days and 16 packs are held as safety stock. What should be the order quantity if the inventory at the time of order placement is 71 packs? Note: answer with the number only. For example, answer like 12
Answer #1

Average daily demand (d) = 17 packs

Review period(R) = 1 week = 7 days

Lead time(L) = 4 days

Safety stock (S) = 16 packs

On hand inventory (I) = 71 packs

Order quantity = d(R+L) + S - I

= 17(7+4) + 16 - 71

= (17×11) + 16 - 71

= 187 + 16 - 71

= 203-71

= 132